OpenStudy (anonymous):

solve the system of equations and give the value of x. {4x-3y=-42.... -6x+7y=68}... a. x=-10. bx=-9. c. x=9. d. x=2

OpenStudy (anonymous):

here u have unique solution

OpenStudy (anonymous):

u here multiply the first eq by 7 and second one by 3 we have and then add 28x-21y=-294 -18x+21y=204 after adding we have 10x = -90 hence x=-9

OpenStudy (anonymous):

ok.. so.. for example; -x+6y=-2... -4x+24y=8 would be 4x-24y=8, -4x+24y=8.. so therefore, no solution? or 0?

OpenStudy (anonymous):

im definitely becoming your fan! :) Thanks for all your help.

OpenStudy (anonymous):

are u multiplying -x+6y=-2 by 4 then it results in -4x+24y=-8

OpenStudy (anonymous):

thanks dear...

OpenStudy (anonymous):

well i mulyiplied by 4.. so should i have done -4?

OpenStudy (anonymous):

so is it x=2? x=0? x=9? or no solution.. ? and how..

OpenStudy (anonymous):

i thought we made the x or y opposite to cancel it out?

OpenStudy (anonymous):

there is -2 at rthe right side of -x+6y=-2

OpenStudy (anonymous):

can u restate the question please

OpenStudy (anonymous):

nevermind.. i got confused, i understand why you did -4 instead of 4..

OpenStudy (anonymous):

so is it no solution?

OpenStudy (anonymous):

if u have the lines as -x+6y=-2 and the other one as -4x+24y=8 then yup no solution as the represent distinct parallel lines

OpenStudy (anonymous):

ok.. thanks sweetie.

OpenStudy (anonymous):

thanx bye dear

OpenStudy (anonymous):

bye :)